### v1

@@ Abstract
Security conferences are important venues at which academics and practitioners share knowledge about new attacks and state-of-the-art defenses. Despite this, researchers have not studied who shares information and about which security topics. To address this, our study characterizes the speakers, sponsors, and topics presented at the most prestigious academic and industry conferences. We collect a longitudinal data set that contains 9,728 abstracts and 1,686 sponsors across 4 academic and 6 industry conferences. There is limited knowledge sharing between industry and academia. Conferences vary significantly in the equality of how talks/authorship is distributed across individuals. The topics of academic and industry abstracts display consistent coverage of techniques within the MITRE ATT&CK framework. Top tier academic conferences, as well as DEFCON and Black Hat, inconsistently address the governance, response and recovery functions of the NIST Cybersecurity Framework. Commercial InfoSec and insurance conferences (RSA, Gartner, Advisen and NetDillgience) cover the framework more consistently. Prevention and detection remain the most common topic of talks, with no clear temporal trend.
keywords:
cybersecurity, science of security, sociology, information sharing, quantitative social science
1 Introduction
Information security knowledge is produced by both academic and industry researchers. Norms across both communities differ in terms of questions like what “good” research looks like, and the role of peer review However, both communities value gathering in dimly lit conference rooms to listen to researchers explain findings with the support of a slide deck.
To attend conferences, attendees travel to foreign cities, sometimes across continents, and spend thousands of dollars on registration. Although some of these activities are not directly related to formal knowledge exchange (e.g. hallway chats and other extracurricular activities), most participants spend hours listening to talks providing knowledge about novel attacks, mitigation measures, lessons learned, and security concepts. For some attendees in both academia and industry, these conferences are akin to a pilgrimage.
Despite the attendees and their employers investing considerable time and money, security conferences have not been studied as part of the security information sharing ecosystem skopik2016problem ; laube2017strategic ; ainslie2023cyber . We are not aware of research into the speakers or topics at industry conferences. Studies of academic security conferences focus on bibliometrics wendzel2020not ; balzarot2022 and which evaluation method was used in a study carver2016establishing ; burcham2017characterizing .
Further investigation is needed because conferences are not a neutral forum to exchange ideas. While conferences provide a platform for selected researchers to present their work, the low acceptance rates mean other researchers are silenced. In selecting the talks that thousands of attendees watch, program committees control which security topics receive attention. This may influence which solutions are built and implemented. Further, paid speaker slots at industry conferences undermine the so-called market for ideas drezner2017ideas .
These potential problems and the lack of prior research motivate a study that characterizes prestigious security conferences. Our study asks: RQ1: Who speaks at security conferences?; RQ2: About which topics?; and RQ3: How does this vary by conference and over time? We collect a sample of titles, abstracts and speakers/authors across $4$ academic and $6$ industry conferences from 2014 to 2022. To analyze topics, we classified the abstracts into offensive, defensive, or neutral (for theoretical papers). We then mapped offensive talks to the MITRE ATT&CK Framework [^1] https://attack.mitre.org/ (accessed 31 July 2023), and defensive talks to the NIST Cybersecurity Framework [^2] https://www.nist.gov/cyberframework (accessed 31 July 2023).
@@ 1 Introduction
The rest of this paper is structured as follows. Section 2 discusses related work in the field of information sharing. Section 3 describes our research methodology. Section 4 presents the results of our study. Section 5 discusses the limitations of the study. Section 6 discusses the implications of our results and makes recommendations. Section 7 concludes the paper.
2 Related Work
@@
Defenders use and share many forms of security information to better calibrate their defensive posture to the threat landscape. Most research focuses on sharing real-time technical indicators laube2017strategic , which we review in Section 2.1. Other communication channels exist by which defenders can learn about security, which are covered in Section 2.2.
2.1 Information Sharing
@@
The two most common forms of security information sharing are threat intelligence and vulnerability management. A range of schema and standards have been introduced to facilitate sharing kampanakis2014security , which includes MAEC, OVAL, XCCDF, CPE, CVE, CVSS and so on. Governments have passed regulations to support information sharing yang2020impact , funded Computer Emergency Response Teams to coordinate vulnerability disclosure slayton2020trusting , and directly shared threat intelligence [^3] https://www.cisa.gov/known-exploited-vulnerabilities-catalog. Commercial providers also offer threat intelligence (TI) and surrounding services wagner2019cyber ; sauerwein2017threat ; brown2015cyber ; sauerwein2019analysis . The associated platforms provide information enrichment and analysis functions to obtain targeted and actionable threat intelligence sauerwein2021threat ; dandurand2013towards ; brown2015cyber ; chismon2015threat .
@@
Recent empirical studies have questioned the quality of threat intelligence. Li et al. li2019reading find little overlap between different TI feeds and cannot identify a way to distinguish quality, which confirms findings from other studies thomas2016abuse ; bouwman2022helping . Expert interviews with TI practitioners reveal processes to evaluate quality are informal bouwman2020different , which is to be expected given this remains an open research challenge zibak2022threat . Encouragingly, a study of threat intelligence during COVID-19 found evidence that collaboration, in the form of aggregating commercial TI feeds, improved coverage of threats bouwman2022helping .
@@
Turning to vulnerability management, bug bounties have proliferated, run both independently by vendors zhao2017devising and on platforms like HackerOne sridhar2021hacking . Empirical evaluations support the efficacy of bug bounty programs over internal testing finifter2013empirical ; zhao2015empirical , even though they follow similar processes votipka2018hackers . However, many firms do not patch even when the vulnerability is known and the patch is available frei2006large ; li2017large . This has motivated research into understanding expert users who apply patches li2019keepers ; mathur2016they , and notifications that nudge firms towards patching stock2016hey ; li2016you ; ccetin2019tell .
@@
So much information is shared about security vulnerabilities that prioritizing what to pay attention to has become a research problem jacobs2020improving ; de2023no . Information overload is also a problem for threat intelligence bouwman2020different . This motivates research exploring online advice and knowledge sharing that may be easier for the recipients to digest than technical indicators.
2.2 Advice and Knowledge Sharing
@@
Online forums are often used to solve security and privacy problems li2023itsup ; tahaei2020understanding , including by developers acar2016you . In some cases, this involves copying insecure code snippets directly from the forum post fischer2017stack . This has motivated research into how to nudge developers towards better choices fischer2019stack .
@@
Online security advice targeted at the general public is also widely available redmiles2016think . An analysis of over a thousand sources of advice reveals that the specific recommendations are actionable and comprehensible, albeit with room for improvement redmiles2020comprehensive . However, the sources recommended 374 unique behaviors, which leads to information overload redmiles2020comprehensive . Some users address the quality issue by looking to trusted sources redmiles2016think .
@@
Conferences address this issue by curating the agenda, via peer review in academia. Research into what information is shared at these conferences has been small-scale. For example, Carver et al. carver2016establishing identify which evaluation method (e.g., proof or empirical) was used by each paper in the 2015 IEEE Security and Privacy proceedings. This study was extended to cover the 2016 ACM CCS proceedings burcham2017characterizing . A bibliometric analysis was conducted of around 30,000 academic computer security papers wendzel2020not , focusing on statistics like the number of words, authors, and citations in each paper.
@@
3 Methodology
We focused on the most prestigious conferences because they have the most influence. For each talk, one could analyze article PDFs, video recordings of talks, or meta information like abstracts and titles. Although full academic articles contain the most information, the industry equivalent (white papers) are not available for all talks. Similarly, video recordings of talks are not consistently available, and automated transcription is imperfect. For these reasons, we analyzed abstracts, which are consistently available across industry and academia.
Section 3.1 describes our corpus of conference abstracts. Section 3.2 explains how we used natural language processing to classify what each talk was about, and Section 3.3 describes how we validated this classification. Section 3.4 discusses ethical considerations and open science. Figure 1 provides a high-level overview of our research design.
Figure 1: Our 2-stage classification uses OpenAI’s gpt model in the first step and then both gpt and cosine similarity to map the abstracts to cybersecurity frameworks. We also collect the sponsors of each conference.
3.1 Data Collection
@@
In academia, we focused on ‘Top Tier’ computer security conferences soneji2022flawed , namely USENIX Security, IEEE S&P, ACM CCS, and NDSS. We did not consider cryptography venues because they do not map onto either framework (see Section 3.2). Journals were excluded because the lack of in-person presentations undermines the comparison to industry conferences.
@@
For industry, we conducted a web search for articles featuring the ‘Top Security Conferences’. Only Black Hat, DEFCON and RSA were consistently included in all of these articles. For example, the SANS Institute runs influential topic-specific conferences (e.g. Threat Intelligence or Network Security), but no specific event is consistently included as the flagship event. We consider these seven conferences to be the most prestigious computer security conferences.
@@
To diversify our sample, we included Gartner Security because of its influence over procurement decisions and also the two major cyber insurance conferences (Advisen and NetDiligence). Insurance is increasingly used by firms to manage cybersecurity risk dambra2020sok . This makes for an interesting comparison because the insurance community’s roots are in finance rather than IT. Table 1 displays the conferences in our sample, along with some simple statistics.
@@
Conference Sample Window Type Talks per year Characters in title Words per abstract Sponsors Gini Index
$\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$ 1 1/N
Black Hat 2014 - 2022 industry (InfoSec) 252 53.83 54.82 26.43 170.98 99.16 179.0 54.29 0.234 0.381
RSA 2017 - 2022 industry (InfoSec) 338.33 34.98 55.69 17.18 75.75 26.98 75.0a 0.0a 0.288 0.441
DEFCON 2014 - 2022 industry (InfoSec) 92.22 32.19 56.10 25.75 176.72 77.72 0 0 0.179 0.382
Gartner Security 2022 industry (InfoSec) 321.0 0.0 68.07 19.38 66.85 20.38 N/A* N/A* 0.328 0.427
Advisen 2014 - 2022 industry (Insurance) 10.0 4.92 38.81 17.62 58.67 22.25 18.13b 5.95b 0.244 0.383
NetDiligence 2019 - 2022 industry (Insurance) 25.0 6.22 51.50 20.35 36.76 27.36 56.50 15.32 0.159 0.269
USENIX Security 2014 - 2022 academic 132.78 68.33 72.30 22.40 184.12 62.10 15.22 3.33 0.284 0.417
IEEE S&P 2014 - 2022 academic 82.22 34.08 69.77 21.42 222.52 63.53 14.67 6.22 0.280 0.380
ACM CCS 2014 - 2022 academic 166.33 34.47 67.99 21.44 186.21 82.81 13.87c 3.33c 0.325 0.451
NDSS 2014 - 2022 academic 73.0 15.80 72.11 19.79 186.54 82.22 10.14d 3.27d 0.209 0.376
Table 1: Descriptive statistic display the mean value across all years. The Gini index is calculated with the total number of talks per individual across all years. a contains data only from year 2022. b year 2020 is missing. c year 2018 is missing. d only years from 2016 - 2022. N/A* when sponsor data was not available.
@@
The data was obtained from the conference websites or extracted directly from the proceedings via web content mining. We used a semi-automated process to extract information about speakers, affiliations, talk titles, abstracts, and dates. The academic talks include keynotes and paper presentations for which an abstract was available, but we did not include poster sessions. For industry conferences, we considered all talks with an abstract listed in the conference agenda and ignored sessions such as networking sessions, welcome remarks, and opening speeches.
The data set comprises a total of 9,728 talks (5,639 industry talks and 4,089 academic talks). We also collected 1,686 entries for sponsors of information security conferences, 451 for academic conferences, and 371 for cyber insurance. Some firms sponsored conferences for multiple years.
3.2 Data Analysis

### v2

@@ Abstract.
Security conferences are important venues for information sharing, where academics and practitioners share knowledge about new attacks and state-of-the-art defenses. Despite their importance, researchers have not systematically examined who shares information and which security topics are discussed. To address this gap, our paper characterizes the speakers, sponsors, and topics presented at prestigious academic and industry security conferences. We compile a longitudinal dataset containing 9,728 abstracts and 1,686 sponsors across four academic and six industry conferences. Our findings show limited information sharing between industry and academia. Conferences vary significantly in how equitably talks and authorship are distributed across individuals. The topics of academic and industry abstracts display consistent coverage of techniques within the MITRE ATT&CK framework. Top-tier academic conferences, as well as DEFCON and Black Hat, address the governance, response, and recovery functions of the NIST Cybersecurity Framework inconsistently. Commercial information security and insurance conferences (RSA, Gartner, Advisen and NetDiligence) more consistently cover the framework. Prevention and detection were the most common topics in the sample period, with no clear temporal trends.
$c$s: Security and privacy Economics of security and privacyccs: Security and privacy Social aspects of security and privacyccs: Social and professional topics Computing industrycopyright: ccjournal: DTRAPjournalyear: 2026journalvolume: 1journalnumber: 1publicationmonth: 1doi: 10.1145/3788676
1. Introduction
Information security knowledge is produced by both academic and industry researchers. However, these two communities differ in their norms and values, including what constitutes “rigorous” research and the role of peer review. Despite these differences, both academic and industry researchers value the exchange of knowledge that occurs when experts convene to present and discuss findings at conferences.
For many professionals in both academia and industry, these events represent important venues for professional development and community engagement. Attendees often travel internationally and spend significant amounts on registration fees and related expenses. Although some of these activities are not directly related to formal knowledge exchange (e.g., hallway discussions and informal extracurricular activities), most participants devote time to listening to presentations that convey insights into novel attacks, mitigation techniques, lessons learned, and emerging security concepts.
Despite the substantial investment of time and resources by attendees and their employers, security conferences have not been systematically studied as part of the security information sharing ecosystem (Skopik et al., 2016; Laube and Böhme, 2017; Ainslie et al., 2023). To our knowledge, no prior research has analyzed the speakers or topics featured at industry conferences. Existing studies on academic security conferences have primarily focused on bibliometrics (Wendzel et al., 2020; Balzarotti, 2022) and evaluation methodologies (Carver et al., 2016; Burcham et al., 2017). This research gap is significant because conferences influence which security topics receive attention, thereby shaping the direction of research and practice. While conferences provide a platform for selected researchers to present their work, low acceptance rates necessarily exclude many others. [^1] The acceptance rates at top-tier academic conferences like IEEE Security and Privacy, USENIX Security, ACM CCS, and NDSS range from 5–20% depending on the year. Industry conferences like Black Hat, DEFCON and RSA do not report acceptance rates.
To address the research gap, our study characterizes information sharing at leading academic and industry security conferences by posing the following research questions:
  • RQ1:
    Who speaks at security conferences?
  • RQ2:
    About which topics?
  • RQ3:
    How does this vary by conference and over time?
We collected a dataset containing titles, abstracts, and author or speaker information from four academic and six industry conferences held between 2014 and 2022. To analyze thematic content, we classified abstracts as offensive, defensive, or neutral (for theoretical work). Offensive talks were mapped to the MITRE ATT&CK Framework [^2] https://attack.mitre.org/ (accessed 19 October 2025), and defensive talks to the NIST Cybersecurity Framework [^3] https://www.nist.gov/cyberframework (accessed 19 October 2025).
Our contributions are as follows:
  • RQ1:
    Among academic conferences, NDSS exhibits the most equitable distribution of speakers, whereas ACM CCS shows the least. Cross-participation between academic, industry, and insurance conferences is rare. We found no evidence that sponsorship significantly affects speaker selection at industry conferences.
  • RQ2:
    Defensive talks are more common than offensive talks at all conferences except DEFCON. Defensive talks focus on protective technology, detection, and response. Offensive talks are more evenly distributed across offensive security categories in MITRE ATT&CK.
  • RQ3:
    The distribution of topics remains relatively stable over time. Compared among conferences, DEFCON covers similar topics to academic conferences. RSA and Gartner emphasize on governance, while cyber insurance conferences focus on incident response.
  • Data:
    The dataset compiled for this study was made publicly available to support replication and further analysis (Walter et al., 2024).
@@ 1. Introduction
The remainder of this paper is structured as follows. Section 2 discusses related work in the field of information sharing. Section 3 describes the research methodology. Section 4 presents the empirical results. Section 5 discusses the limitations of the study. Section 6 outlines the implications of our findings and provides recommendations. Section 7 concludes the paper.
2. Related Work
@@
Defenders use and share various forms of security information to better calibrate their defensive posture to the threat landscape (Webb et al., 2014; Ahmad et al., 2021). Most research focuses on sharing real-time technical indicators (Laube and Böhme, 2017), which we review in Section 2.1. Other communication channels through which defenders acquire and disseminate security knowledge are discussed in Section 2.2.
2.1. Information Sharing
@@
The two most common forms of security information sharing are threat intelligence (TI) and vulnerability management. A range of schema and standards have been introduced to facilitate sharing (Kampanakis, 2014), including MAEC, OVAL, XCCDF, CPE, CVE, and CVSS. Governments have passed regulations to support information sharing (Yang et al., 2020), funded Computer Emergency Response Teams (CERTs) to coordinate vulnerability disclosure (Slayton and Clarke, 2020), and directly shared threat intelligence [^4] https://www.cisa.gov/known-exploited-vulnerabilities-catalog (accessed 19 October 2025). Commercial providers also offer threat intelligence and surrounding services (Wagner et al., 2019; Sauerwein et al., 2017; Brown et al., 2015; Sauerwein et al., 2019). The associated platforms provide information enrichment and analysis functions to obtain targeted and actionable threat intelligence (Sauerwein et al., 2021; Dandurand and Serrano, 2013; Brown et al., 2015; Chismon and Ruks, 2015).
@@
Recent empirical studies have questioned the quality of threat intelligence. Li et al. (Li et al., 2019b) find little overlap between different TI feeds and cannot identify a way to distinguish quality, which confirms findings from other studies (Thomas et al., 2016; Bouwman et al., 2022). Expert interviews with TI practitioners reveal processes to evaluate quality are informal (Bouwman et al., 2020), which is to be expected given this remains an open research challenge (Zibak et al., 2022). Encouragingly, a study of threat intelligence during COVID-19 found evidence that collaboration, in the form of aggregating commercial TI feeds, improved coverage of threats (Bouwman et al., 2022).
@@
In the area of vulnerability management, bug bounty programs have proliferated, run both independently by vendors (Zhao et al., 2017) and on platforms like HackerOne (Sridhar and Ng, 2021). Empirical evaluations support the efficacy of bug bounty programs over internal testing (Finifter et al., 2013; Zhao et al., 2015), even though they follow similar processes (Votipka et al., 2018). Recent work has shown bug hunters in China organize into teams, in part to facilitate information sharing (Piao et al., 2025). However, many firms do not patch even when the vulnerability is known and the patch is available (Frei et al., 2006; Li and Paxson, 2017). This has motivated research into understanding expert users who apply patches (Li et al., 2019a; Mathur et al., 2016), and notifications that nudge firms towards patching (Stock et al., 2016; Li et al., 2016; Çetin et al., 2019).
@@
The sheer volume of information shared about security vulnerabilities has created a prioritization challenge, making it difficult for defenders to identify which issues merit attention (Jacobs et al., 2020; de Smale et al., 2023). Information overload similarly affects threat intelligence (Bouwman et al., 2020). This motivates research exploring online advice and knowledge sharing that may be easier for the recipients to digest than technical indicators.
2.2. Advice and Knowledge Sharing
@@
Online forums are often used to solve security and privacy problems (Li et al., 2023; Tahaei et al., 2020), including by developers (Acar et al., 2016). In some cases, this involves copying insecure code snippets directly from the forum post (Fischer et al., 2017). This has motivated research into how to nudge developers towards better choices (Fischer et al., 2019).
@@
Online security advice targeted at the general public is also widely available (Redmiles et al., 2016). An analysis of over a thousand sources of advice reveals that the specific recommendations are actionable and comprehensible, albeit with room for improvement (Redmiles et al., 2020). However, the sources recommended 374 unique behaviors, which leads to information overload (Redmiles et al., 2020). Some users address the quality issue by looking to trusted sources (Redmiles et al., 2016).
@@
Conferences address this issue by curating the agenda, via peer review in academia. Research into what information is shared at these conferences has been small-scale. For example, Carver et al. (Carver et al., 2016) identify which evaluation method (e.g., proof or empirical) was used by each paper in the 2015 IEEE Security and Privacy proceedings. This study was extended to cover the 2016 ACM CCS proceedings (Burcham et al., 2017). A bibliometric analysis was conducted of around 30,000 academic computer security papers (Wendzel et al., 2020), focusing on statistics like the number of words, authors, and citations in each paper.
@@
3. Methodology
We focused on prestigious security conferences, as they exerted the greatest influence within both academia and industry. Each talk can be analyzed as a video recording, via the associated paper, or metadata such as the abstract and title. While full academic papers contained the most comprehensive information, their industry equivalents (e.g., white papers) were not consistently available. Similarly, video recordings of talks were inconsistently published, and automated transcription remained imperfect. For these reasons, we relied on abstracts and titles, which were consistently available across both domains.
Section 3.1 describes our corpus of conference abstracts. Section 3.2 explains the natural language processing (NLP) techniques used to classify talks. Section 3.3 outlines our validation strategy. Section 3.4 discusses ethical considerations and open science. Figure 1 provides a high-level overview of our research design.
Figure 1. Two-stage classification process. In the first stage, GPT-based natural language processing is used to classify abstracts. In the second stage, both GPT and cosine similarity are applied to map abstracts to cybersecurity frameworks.
3.1. Data Collection
@@
For academia, we focused on top-tier computer security conferences (Soneji et al., 2022): USENIX Security, IEEE S&P, ACM CCS, and NDSS. Cryptography-specific venues were excluded because their focus did not align with the frameworks considered (see Section 3.2). Journals were also excluded, as the absence of in-person presentations undermines the comparability with industry conferences.
@@
For industry, we conducted a structured web search for articles featuring the ‘top security conferences’. Only Black Hat, DEFCON and RSA were consistently included across all identified lists. For example, the SANS Institute hosts influential topic-specific conferences (e.g. Threat Intelligence or Network Security), but no specific event is consistently included as the flagship event. We consider these seven conferences to be the most prestigious computer security conferences.
@@
To diversify our sample, we included Gartner Security - given its influence over procurement decisions - and two leading cyber insurance conferences (Advisen and NetDiligence). Cyber insurance is increasingly used by firms to manage cybersecurity risk (Dambra et al., 2020). Its inclusion offers an informative comparison, particularly because the insurance community’s roots lay in finance rather than IT. Table 1 displays the conferences in our sample, along with some simple statistics.
@@
Conference Sample Window Type Talks per year Characters in title Words per abstract
$\mu$ $\sigma$ $\mu$ $\sigma$ $\mu$ $\sigma$
Black Hat 2014 - 2022 industry (InfoSec) 252 53.83 54.82 26.43 170.98 99.16
RSA 2017 - 2022 industry (InfoSec) 338.33 34.98 55.69 17.18 75.75 26.98
DEFCON 2014 - 2022 industry (InfoSec) 92.22 32.19 56.10 25.75 176.72 77.72
Gartner Security 2022 industry (InfoSec) 321.0 0.0 68.07 19.38 66.85 20.38
Advisen 2014 - 2022 industry (Insurance) 10.0 4.92 38.81 17.62 58.67 22.25
NetDiligence 2019 - 2022 industry (Insurance) 25.0 6.22 51.50 20.35 36.76 27.36
USENIX Security 2014 - 2022 academic 132.78 68.33 72.30 22.40 184.12 62.10
IEEE S&P 2014 - 2022 academic 82.22 34.08 69.77 21.42 222.52 63.53
ACM CCS 2014 - 2022 academic 166.33 34.47 67.99 21.44 186.21 82.81
NDSS 2014 - 2022 academic 73.0 15.80 72.11 19.79 186.54 82.22
Table 1. Conference metadata. Descriptive statistics showing the mean ($\mu$) and standard deviation ($\sigma$) across all years for the number of talks per year, title length in characters, and abstract length.
Conference Sponsors
$\mu$ $\sigma$ Notes
Black Hat 179.0 54.29
RSA 75.0 0.0 contains data only from year 2022.
DEFCON 0 0
Gartner Security N/A N/A sponsor data was not available
Advisen 18.13 5.95 year 2020 is missing.
NetDiligence 56.50 15.32
USENIX Security 15.22 3.33
IEEE S&P 14.67 6.22
ACM CCS 13.87 3.33 year 2018 is missing.
NDSS 10.14 3.27 only years from 2016 - 2022
Table 2. Sponsorship data. Mean ($\mu$) and standard deviation ($\sigma$) of the number of sponsors per conference. Notes indicate data-specific limitations.
Conference Gini Index
1 1/N
Black Hat 0.234 0.381
RSA 0.288 0.441
DEFCON 0.179 0.382
Gartner Security 0.328 0.427
Advisen/Zywave 0.244 0.383
NetDiligence 0.159 0.269
USENIX Security 0.284 0.417
IEEE S&P 0.280 0.380
ACM CCS 0.325 0.451
NDSS 0.209 0.376
Table 3. Speaker distribution equality. The Gini index is calculated for the total number of talks per individual across all years. The two columns report Gini index values under different normalization factors (1 and 1/N).
@@
The dataset used in this study was collected exclusively from publicly available sources, without reliance on institutional subscriptions or paywalled materials. Data was retrieved directly from official conference websites or openly accessible conference proceedings using web content mining techniques. We extracted freely available metadata, including titles, author names, abstracts, and sponsor lists, as published on conference or publisher landing pages. Consequently, no special permissions were required from conference organizers or authors. A semi-automated process was employed to extract information on speakers, affiliations, titles, abstracts, and dates. For academic venues, we included keynotes and paper presentations for which an abstract was available, excluding poster sessions. For industry conferences, we considered all talks with an abstract listed in the conference agenda and omitted non-technical sessions such as networking events or welcome remarks. The resulting dataset comprises 9,728 talks (5,639 industry; 4,089 academic) and 1,686 sponsor entries, of which 451 corresponded to academic and 371 to cyber insurance conferences.
3.2. Data Analysis